A certain weak acid, HA, with a Ka value of 5.61×10−6, is titrated with NaOH.More strong base is added until the equivalence point is reached. What is the pH of this solution at the equivalence point if the total volume is 30.0 mL ?
1
Expert's answer
2018-07-18T04:42:58-0400
If we assume that the concentrations of acid and base are equal [base] = [acid], then the Henderson-Hasselbalch equation:

pH = pKa + log ([base]/[acid])

[base] = molar concentration of a conjugate base

[acid] = molar concentration of a undissociated weak acid (M)

Ka = 5.61×10^−6 pKa = 5.25

pH = 5.25